OpenStudy (anonymous):

help. cube root of -2i in r (costheta +isintheta) form?

OpenStudy (trojanpoem):

let z = -2i x = 0 , y = -2 theta = -2/0 , theta = 270 |r| = sqrt( 0 + 4) = 2 z = |r| (costheta + sintheta) = 2 ( cos270 + i(sin270)) You can easily verify it, by calculating the value of sin , cos, multiplying times the 2 out.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

how would I write it with cuberoot(2) as my r value?

OpenStudy (irishboy123):

you should have 3 cube roots

OpenStudy (anonymous):

nevermind, I think I figured it out, since cbrt(2) = 2^(1/3), all I do is divide 270 by 3 and get cbrt(2)(cos(90)+isin(90))

OpenStudy (anonymous):

or not, it appears it was 210 :(

OpenStudy (irishboy123):

 \(\large -2i = 2 e^{i \frac{3 \pi}{2}}\) but also \(\large -2i = 2 e^{i (\frac{3 \pi}{2}+2 \pi)}\), \(\large -2i = 2 e^{i (\frac{3 \pi}{2}+4 \pi)}\) , ...... so \(\large -2i = 2e^{i\frac{3 \pi}{2}}, 2e^{i\frac{7 \pi}{2}} , 2e^{i\frac{11 \pi}{2}} , 2e^{i\frac{15 \pi}{2}},....\) so \(\sqrt[3]{-2i } = \sqrt[3]{2}e^{i\frac{ \pi}{2}}, \sqrt[3]{2}e^{i\frac{7 \pi}{6}} , \sqrt[3]{2}e^{i\frac{11 \pi}{6}} , \sqrt[3]{2}e^{i\frac{15 \pi}{6}},....\) so the fourth root is a repeat of the first as \( \frac{15 \pi}{6} = \frac{\pi }{2} + 2 \pi \); and they repeat ever after.... so in cis form you have \(\sqrt[3]{-2i } = \sqrt[3]{2}\; cis \, 90, \sqrt[3]{2} \; cis\, 210 , \sqrt[3]{2} \; cis \,330 ,\)

OpenStudy (irishboy123):

but the 90deg one would be the *principal* cube root of you are looking for only one answer....